# ul. Pasteura 5, 02-093 Warsaw, Poland

### v1

@@ 3.1.2 Matter domination from moduli decay(s)
In the usual picture, moduli can generically be displaced from the minimum of their potential after inflation. This is especially for the volume modulus, [^22] The same is true for the dilaton. The attentive reader may worry that displaced, oscillating moduli would quickly overtake the energy density of the kinating volume after inflation. However, such moduli do not redshift as matter if their mass is decreasing with time. In LVS, for example, the energy density of the complex structure moduli (including the axio-dilaton) redshifts exactly as $\rho^{-6}$ during kination Apers:2024dtn . which couples to all energy sources that are present (for example, the inflaton potential). If the displacement is small, when the Hubble rate drops down to $H=\mathcal{O}\left(m_{\Phi}\right)$ it will start to oscillate, giving rise to an early epoch of matter domination lasting until the moduli decay. To distinguish it from other scenarios of early matter domination discussed in the literature, we shall denote this epoch as moduli-induced Early Matter Domination (mEMD).
@@
$t_{m}\simeq{\rm Max}\left\{\frac{1}{m_{\Phi}},\frac{1}{H(t_{{\rm min}})}\right\},$ (38)
where $t_{{\rm min}}$ is the time it takes to reach the minimum. In the specific case of LVS, however, it turns out that the two coincide, as $t_{m}\sim\frac{1}{m_{\Phi}}\sim\frac{1}{H(t_{{\rm min}})}$. To see this, let us refer back to Sec. 3.1.1 and notice that the tracker regime has to be attained before $\Phi$ reaches the minimum, in order to avoid the overshoot problem. One can then use the fact that $y$ is approximately constant (i.e. (37)) to relate the volume to the Hubble scale, resulting in
$\frac{H(t_{{\rm min}})}{M_{P}}=\frac{\lambda}{3}\left(\frac{2V_{0}}{(2-\gamma)\gamma M_{P}^{4}}\right)^{1/2}e^{-\frac{\lambda(\Phi_{{\rm fin}}-\Phi_{0})}{2M_{P}}}.$ (39)
@@
$m_{\Phi}\sim\sqrt{\frac{|V(\Phi_{{\rm fin}})|}{M_{P}^{2}}}\sim\sqrt{\frac{V_{0}}{M_{P}^{2}}}e^{-\frac{\lambda(\Phi_{{\rm fin}}-\Phi_{0})}{2M_{P}}},$ (40)
which aside from $\mathcal{O}(1)$ factors has the exact same dependence on the field as Eq. (39). Therefore, at least in this scenario one can never really disentangle the two scales $H(t_{{\rm min}})$ and $m_{\Phi}$ in a parametric manner.
@@
$\Gamma_{\Phi_{i}}=\frac{c_{i}}{48\pi}\frac{m^{3}_{\Phi}}{M_{P}^{2}},$ (41)
with $c_{i}$ a model-dependent, $\mathcal{O}(1)$ factor. Therefore, the last modulus to decay is always the lightest, which in the scenarios we are considering is typically the volume $\Phi$. In the approximation of instantaneous decay, valid as long as $\Gamma_{\Phi}\gg H$, the universe will be reheated at a time when $3H\simeq\Gamma_{\Phi}$, corresponding to a reheating temperature of Cicoli:2012aq ; Hebecker:2014gka
$T_{\text{reheat }}\sim\sqrt{\Gamma_{\Phi}M_{P}}\sim\mathcal{O}(1)\mathrm{GeV}\left(\frac{m_{\Phi}}{3\times 10^{6}\mathrm{GeV}}\right)^{3/2}.$ (42)
To be compatible with cosmological observations, $T_{\text{reheat }}$ must be well above the MeV scale Coughlan:1983ci ; Banks:1993en ; deCarlos:1993wie , resulting in the bound $m_{\Phi}\gtrsim 30\,{\rm TeV}$, which is surprisingly saturated by typical values of microscopic parameters for cases of interest in Sec. 5. Moreover, volume decays can contribute to the dark radiation problem, and are constrained by bounds on the effective number of neutrino species $N_{eff}$ Planck:2018vyg . In LVS this is a significant problem, as there is always one decay channel (the volume axion) with an $\mathcal{O}(1)$ branching ratio Cicoli:2012aq ; Higaki:2012ar ; Hebecker:2014gka (see also Leedom:2024qgr for possible enhancements due to parametric resonance).
A generic solution was recently suggested in Cicoli:2022fzy , by realizing that in models with a high supersymmetry-breaking scale the Higgs inherits a parametrically enhanced coupling to the volume modulus due to the fine tuning of its mass. The idea is that the trilinear coupling between the Higgs and volume should not be mediated by the (observed) fine-tuned Higgs mass $m_{H}$, but rather feel the effect of its un-tuned value, whose magnitude is set by the soft terms (and hence proportional to the gravitino mass $m_{3/2}$). In formulae, the unnaturally small Higgs mass arises from the highly fine-tuned cancellation between the tree-level and one-loop term at the stabilised value for the volume: [^25] In (43), $\alpha_{0}$ and $\alpha_{\rm{loop}}$ are some generic tree-level and one-loop coefficients, $m_{KK}$ the Kaluza-Klein scale and $W_{0}$ the flux contribution to the superpotential.
$m_{H}^{2}\sim m_{3/2}^{2}\left[\alpha_{0}+\alpha_{\mathrm{loop}}\ln\left(\frac{m_{\mathrm{KK}}}{m_{3/2}}\right)\right]\simeq\left(\frac{W_{0}}{\mathcal{V}}\right)^{2}\left[\alpha_{0}+\alpha_{\text{loop }}\ln\left(\frac{\mathcal{V}^{1/2}}{W_{0}}\right)\right]M_{P}^{2}.$ (43)
However, the volume dependence of the two terms is different, and expanding $\mathcal{V}$ in terms of the canonically normalized volume (22) around $\mathcal{V}=e^{\sqrt{\frac{3}{2}}\frac{\Phi_{0}+\delta\Phi}{M_{P}}}$ leads to a trilinear coupling of the form
$\mathcal{L}\supset\frac{1}{2}\sqrt{\frac{3}{2}}\alpha_{\rm{loop}}m_{3/2}^{2}h^{2}\frac{\delta\Phi}{M_{P}}\sim m_{3/2}^{2}\alpha_{\text{loop }}h^{2}\frac{\delta\Phi}{M_{P}},$ (44)
@@
$\Gamma_{\Phi\rightarrow hh}\simeq\frac{1}{16\pi}\frac{m_{3/2}^{4}\alpha_{\rm{loop}}^{2}}{m_{\Phi}M_{P}^{2}}\sim\left(\alpha_{\rm{loop}}\mathcal{V}\right)^{2}\frac{m_{\Phi}^{3}}{M_{P}^{2}},$ (45)
which is parametrically enhanced with respect to (41), and thus dominates over all other decay channels. If the lifetime of the volume receives such a large suppression, it is however possible for heavier moduli, whose decay rate is still of the form (41), to outlive it. This will depend on the fine details of the model, see Cicoli:2022fzy for some examples. Another caveat is that, in the above, we have implicitly assumed that the soft terms are of the same order as the gravitino mass, $m_{\rm{soft}}\sim m_{3/2}$, the naive expectation for gravity-mediated SUSY breaking. This is not true for sequestered scenarios where the SM brane is localized on a small cycle away from the bulk of the Calabi-Yau, where soft terms can be hierarchically smaller thanks to certain cancellations Blumenhagen:2009gk ; Aparicio:2014wxa (see also Reece:2015qbf for an EFT analysis). In particular, depending on the precise model building details (for example the uplifting mechanism) one can have soft scalar masses with a parametric scaling
$m_{H}\sim m_{3/2}\sqrt{\frac{m_{3/2}}{M_{P}}}\sim M_{P}\left(\frac{M_{P}}{\mathcal{V}}\right)^{3/2}\quad\quad{\rm{or}}\quad\quad m_{H}\sim\frac{m_{3/2}^{2}}{M_{P}}\sim M_{P}\left(\frac{M_{P}}{\mathcal{V}}\right)^{2}.$ (46)
In either case, the decay rates $\Gamma_{\phi_{b}\rightarrow hh}$ would have no parametric enhancement with respect to the decay rate (41) into volume axions.
@@
$\frac{H_{\rm{m}}}{H_{\rm{reh}}}\simeq\frac{m_{\Phi}}{\Gamma_{\Phi\rightarrow hh}}\simeq\frac{M_{P}^{2}}{m_{\Phi}^{2}}\frac{1}{\left(\alpha_{\rm{loop}}\mathcal{V}\right)^{2}}.$ (47)
The other case is when the last modulus to decay (which could be the volume in the case of suppressed soft terms, but possibly something else) does so with the rate (41), so that
$\frac{H_{\rm{m}}}{H_{\rm{reh}}}\simeq\frac{m_{\Phi}}{\Gamma_{\Phi\rightarrow hh}}\simeq\frac{M_{P}^{2}}{m_{\Phi}^{2}}.$ (48)
@@
  • Vanilla LVS. We assume that the volume modulus decays gravitationally as in Eq. (41) and ignore dark radiation issues. Fitting to the PTA signal, we will see that the vanilla LVS predicts a change of slope in the LISA band, with the particular frequency depending on the mass of the modulus. Due to the gravitational decay, the period of early matter domination is large and dilutes any high-frequency signal.
@@
  • Short Volume Lifetime (SVL-LVS). We consider instead decay as in Eq. (45). The prediction at the LISA band is then as in standard cosmology. However, we will observe that for a large hierarchy between the inflation and late-time scales, the high-frequency signal can be very large with a spectral index that cannot be reproduced in scenarios with constant tension.
@@
Finally, it should be noted that non-standard cosmologies have the potential to boost GW signals other than the ones coming from cosmic superstrings. This is most relevant for sources that are active at very early times, such as the primordial GW background from inflation. In Co:2021lkc ; Gouttenoire:2021wzu ; Gouttenoire:2021jhk (see also Chowdhury:2022gdc ; Chowdhury:2023opo ), it was found that a long, early kination epoch can enhance the primordial GW background to observable levels, and one might worry this could dominate the signal we describe. For the scenarios we consider, however, the spectrum from cosmic superstrings will always be larger by many orders of magnitude. [^26] Best-case scenarios for the GW background from inflation estimate $h^{2}\Omega_{\rm{inf}}\sim 10^{-16}$, around 8 orders of magnitude below the radiation plateau in our benchmark in Section 5. It is a well known fact (and a simple conclusion from Section 4) that the dependence on $G\mu$ of the amplitude of the radiation plateau is $h^{2}\Omega_{\rm{GW}}\sim\sqrt{G\mu}$. For the amplitudes to be equal we would thus require (with probabilities as in our benchmark in Section 5) $G\mu\sim 10^{-28}$, or an internal volume $\mathcal{V}\sim 10^{26}$, which yields a TeV string scale. At low frequencies, where the effect of such modified epochs cannot be felt, the magnitude of $h^{2}\Omega_{\rm{GW}}$ from cosmic superstrings will lie close to the region probed by PTA observations for realistic values of the parameters (such as the string tension), far above the prediction for the inflationary background. As the frequency increases, the amplitude for both signals can either decrease during matter domination (with the same rate) or increase during kination (with a larger rate for superstrings). [^27] Only in the case of varying tension. Also, the turning point frequencies may be shifted. In particular, the effect of the modified expansion rate translates into a peak for the inflationary background, given by
$h^{2}\Omega_{\rm{inf}}^{(\rm{peak})}\simeq h^{2}\Omega_{\rm{inf}}(f_{\Delta})e^{2N_{kin}-N_{m}}.$ (49)
@@
In the above, $f_{\Delta}$ is the frequency corresponding to the end of the matter domination at $t_{\Delta}$, and $N_{kin},N_{m}$ are the e-folds of kination and matter domination respectively. This must compared with (101), which is always larger by a factor $\left(\frac{\mu(t_{0})}{\mu{(t_{\Delta})}}\right)^{2}$, the ratio of the string tensions at the initial and final times.
@@ 3.2 Effect on the string tension
To close of this section, let us examine how the fundamental string tension depends on time in the examples we have considered. Let us remind the reader that this can be obtained by simply plugging in the time-dependence of the compactification volume [^28] And also of the dilaton, which here is constant. in (1). In the case of volume kination, one can combine Eqs (22) and (27) to obtain
$\mu_{\rm{kin}}(t)=\frac{g_{s}^{2}M_{P}^{2}}{4\pi\mathcal{V}_{0}^{\rm{kin}}}\frac{t_{0}^{\rm{kin}}}{t},$ (50)
@@
$\mu_{\rm{tr}}(t)=\frac{g_{s}^{2}M_{P}^{2}}{4\pi\mathcal{V}_{0}^{\rm{tr}}}\left(\frac{t_{0}^{\rm{tr}}}{t}\right)^{\frac{2}{\lambda}\sqrt{\frac{3}{2}}}.$ (51)
@@
$\mu^{LVS}_{\rm{tr}}(t)=\frac{g_{s}^{2}M_{P}^{2}}{4\pi\mathcal{V}_{0}^{\rm{tr}}}\left(\frac{t_{0}^{\rm{tr}}}{t}\right)^{\frac{2}{3}}.$ (52)

### v2

@@ 3.1.2 Matter domination from moduli decay(s)
In the usual picture, moduli can generically be displaced from the minimum of their potential after inflation. This is especially for the volume modulus, [^24] The same is true for the dilaton. The attentive reader may worry that displaced, oscillating moduli would quickly overtake the energy density of the kinating volume after inflation. However, such moduli do not redshift as matter if their mass is decreasing with time. In LVS, for example, the energy density of the complex structure moduli (including the axio-dilaton) redshifts exactly as $\rho^{-6}$ during kination [105]. which couples to all energy sources that are present (for example, the inflaton potential). If the displacement is small, when the Hubble rate drops down to $H=\mathcal{O}\left(m_{\Phi}\right)$ it will start to oscillate, giving rise to an early epoch of matter domination lasting until the moduli decay. To distinguish it from other scenarios of early matter domination discussed in the literature, we shall denote this epoch as moduli-induced Early Matter Domination (mEMD).
@@
$t_{m}\simeq{\rm Max}\left\{\frac{1}{m_{\Phi}},\frac{1}{H(t_{{\rm min}})}\right\},$ (3.17)
where $t_{{\rm min}}$ is the time it takes to reach the minimum. In the specific case of LVS, however, it turns out that the two coincide, as $t_{m}\sim\frac{1}{m_{\Phi}}\sim\frac{1}{H(t_{{\rm min}})}$. To see this, let us refer back to Sec. 3.1.1 and notice that the tracker regime has to be attained before $\Phi$ reaches the minimum, in order to avoid the overshoot problem. One can then use the fact that $y$ is approximately constant (i.e. (3.16)) to relate the volume to the Hubble scale, resulting in
$\frac{H(t_{{\rm min}})}{M_{P}}=\frac{\lambda}{3}\left(\frac{2V_{0}}{(2-\gamma)\gamma M_{P}^{4}}\right)^{1/2}e^{-\frac{\lambda(\Phi_{{\rm fin}}-\Phi_{0})}{2M_{P}}}.$ (3.18)
@@
$m_{\Phi}\sim\sqrt{\frac{|V(\Phi_{{\rm fin}})|}{M_{P}^{2}}}\sim\sqrt{\frac{V_{0}}{M_{P}^{2}}}e^{-\frac{\lambda(\Phi_{{\rm fin}}-\Phi_{0})}{2M_{P}}},$ (3.19)
which aside from $\mathcal{O}(1)$ factors has the exact same dependence on the field as Eq. (3.18). Therefore, at least in this scenario one can never really disentangle the two scales $H(t_{{\rm min}})$ and $m_{\Phi}$ in a parametric manner.
@@
$\Gamma_{\Phi_{i}}=\frac{c_{i}}{48\pi}\frac{m^{3}_{\Phi}}{M_{P}^{2}},$ (3.20)
with $c_{i}$ a model-dependent, $\mathcal{O}(1)$ factor. Therefore, the last modulus to decay is always the lightest, which in the scenarios we are considering is typically the volume $\Phi$. In the approximation of instantaneous decay, valid as long as $\Gamma_{\Phi}\gg H$, the universe will be reheated at a time when $3H\simeq\Gamma_{\Phi}$, corresponding to a reheating temperature of [112, 113]
$T_{\text{reheat }}\sim\sqrt{\Gamma_{\Phi}M_{P}}\sim\mathcal{O}(1)\mathrm{GeV}\left(\frac{m_{\Phi}}{3\times 10^{6}\mathrm{GeV}}\right)^{3/2}.$ (3.21)
To be compatible with cosmological observations, $T_{\text{reheat }}$ must be well above the MeV scale [142, 143, 144], resulting in the bound $m_{\Phi}\gtrsim 30\,{\rm TeV}$, which is surprisingly saturated by typical values of microscopic parameters for cases of interest in Sec. 5. Moreover, volume decays can contribute to the dark radiation problem, and are constrained by bounds on the effective number of neutrino species $N_{eff}$ [126]. In LVS this is a significant problem, as there is always one decay channel (the volume axion) with an $\mathcal{O}(1)$ branching ratio [112, 145, 113] (see also [146] for possible enhancements due to parametric resonance).
A generic solution was recently suggested in [147], by realizing that in models with a high supersymmetry-breaking scale the Higgs inherits a parametrically enhanced coupling to the volume modulus due to the fine tuning of its mass. The idea is that the trilinear coupling between the Higgs and volume should not be mediated by the (observed) fine-tuned Higgs mass $m_{H}$, but rather feel the effect of its un-tuned value, whose magnitude is set by the soft terms (and hence proportional to the gravitino mass $m_{3/2}$). In formulae, the unnaturally small Higgs mass arises from the highly fine-tuned cancellation between the tree-level and one-loop term at the stabilised value for the volume: [^27] In (3.22), $\alpha_{0}$ and $\alpha_{\rm{loop}}$ are some generic tree-level and one-loop coefficients, $m_{KK}$ the Kaluza-Klein scale and $W_{0}$ the flux contribution to the superpotential.
$m_{H}^{2}\sim m_{3/2}^{2}\left[\alpha_{0}+\alpha_{\mathrm{loop}}\ln\left(\frac{m_{\mathrm{KK}}}{m_{3/2}}\right)\right]\simeq\left(\frac{W_{0}}{\mathcal{V}}\right)^{2}\left[\alpha_{0}+\alpha_{\text{loop }}\ln\left(\frac{\mathcal{V}^{1/2}}{W_{0}}\right)\right]M_{P}^{2}.$ (3.22)
However, the volume dependence of the two terms is different, and expanding $\mathcal{V}$ in terms of the canonically normalized volume (3.1) around $\mathcal{V}=e^{\sqrt{\frac{3}{2}}\frac{\Phi_{0}+\delta\Phi}{M_{P}}}$ leads to a trilinear coupling of the form
$\mathcal{L}\supset\frac{1}{2}\sqrt{\frac{3}{2}}\alpha_{\rm{loop}}m_{3/2}^{2}h^{2}\frac{\delta\Phi}{M_{P}}\sim m_{3/2}^{2}\alpha_{\text{loop }}h^{2}\frac{\delta\Phi}{M_{P}},$ (3.23)
@@
$\Gamma_{\Phi\rightarrow hh}\simeq\frac{1}{16\pi}\frac{m_{3/2}^{4}\alpha_{\rm{loop}}^{2}}{m_{\Phi}M_{P}^{2}}\sim\left(\alpha_{\rm{loop}}\mathcal{V}\right)^{2}\frac{m_{\Phi}^{3}}{M_{P}^{2}},$ (3.24)
which is parametrically enhanced with respect to (3.20), and thus dominates over all other decay channels. If the lifetime of the volume receives such a large suppression, it is however possible for heavier moduli, whose decay rate is still of the form (3.20), to outlive it. This will depend on the fine details of the model, see [147] for some examples. Another caveat is that, in the above, we have implicitly assumed that the soft terms are of the same order as the gravitino mass, $m_{\rm{soft}}\sim m_{3/2}$, the naive expectation for gravity-mediated SUSY breaking. This is not true for sequestered scenarios where the SM brane is localized on a small cycle away from the bulk of the Calabi-Yau, where soft terms can be hierarchically smaller thanks to certain cancellations [110, 148] (see also [149] for an EFT analysis). In particular, depending on the precise model building details (for example the uplifting mechanism) one can have soft scalar masses with a parametric scaling
$m_{H}\sim m_{3/2}\sqrt{\frac{m_{3/2}}{M_{P}}}\sim M_{P}\left(\frac{M_{P}}{\mathcal{V}}\right)^{3/2}\quad\quad{\rm{or}}\quad\quad m_{H}\sim\frac{m_{3/2}^{2}}{M_{P}}\sim M_{P}\left(\frac{M_{P}}{\mathcal{V}}\right)^{2}.$ (3.25)
In either case, the decay rates $\Gamma_{\phi_{b}\rightarrow hh}$ would have no parametric enhancement with respect to the decay rate (3.20) into volume axions.
@@
$\frac{H_{\rm{m}}}{H_{\rm{reh}}}\simeq\frac{m_{\Phi}}{\Gamma_{\Phi\rightarrow hh}}\simeq\frac{M_{P}^{2}}{m_{\Phi}^{2}}\frac{1}{\left(\alpha_{\rm{loop}}\mathcal{V}\right)^{2}}.$ (3.26)
The other case is when the last modulus to decay (which could be the volume in the case of suppressed soft terms, but possibly something else) does so with the rate (3.20), so that
$\frac{H_{\rm{m}}}{H_{\rm{reh}}}\simeq\frac{m_{\Phi}}{\Gamma_{\Phi\rightarrow hh}}\simeq\frac{M_{P}^{2}}{m_{\Phi}^{2}}.$ (3.27)
@@
  • Vanilla LVS. We assume that the volume modulus decays gravitationally as in Eq. (3.20) and ignore dark radiation issues. Fitting to the PTA signal, we will see that the vanilla LVS predicts a change of slope in the LISA band, with the particular frequency depending on the mass of the modulus. Due to the gravitational decay, the period of early matter domination is large and dilutes any high-frequency signal.
@@
  • Short Volume Lifetime (SVL-LVS). We consider instead decay as in Eq. (3.24). The prediction at the LISA band is then as in standard cosmology. However, we will observe that for a large hierarchy between the inflation and late-time scales, the high-frequency signal can be very large with a spectral index that cannot be reproduced in scenarios with constant tension.
@@
Finally, it should be noted that non-standard cosmologies have the potential to boost GW signals other than the ones coming from cosmic superstrings. This is most relevant for sources that are active at very early times, such as the primordial GW background from inflation. In [150, 151, 152] (see also [153, 154]), it was found that a long, early kination epoch can enhance the primordial GW background to observable levels, and one might worry this could dominate the signal we describe. For the scenarios we consider, however, the spectrum from cosmic superstrings will always be larger by many orders of magnitude. [^28] Best-case scenarios for the GW background from inflation estimate $h^{2}\Omega_{\rm{inf}}\sim 10^{-16}$, around 8 orders of magnitude below the radiation plateau in our benchmark in Section 5. It is a well known fact (and a simple conclusion from Section 4) that the dependence on $G\mu$ of the amplitude of the radiation plateau is $h^{2}\Omega_{\rm{GW}}\sim\sqrt{G\mu}$. For the amplitudes to be equal we would thus require (with probabilities as in our benchmark in Section 5) $G\mu\sim 10^{-28}$, or an internal volume $\mathcal{V}\sim 10^{26}$, which yields a TeV string scale. At low frequencies, where the effect of such modified epochs cannot be felt, the magnitude of $h^{2}\Omega_{\rm{GW}}$ from cosmic superstrings will lie close to the region probed by PTA observations for realistic values of the parameters (such as the string tension), far above the prediction for the inflationary background. As the frequency increases, the amplitude for both signals can either decrease during matter domination (with the same rate) or increase during kination (with a larger rate for superstrings). [^29] Only in the case of varying tension. Also, the turning point frequencies may be shifted. In particular, the effect of the modified expansion rate translates into a peak for the inflationary background, given by
$h^{2}\Omega_{\rm{inf}}^{(\rm{peak})}\simeq h^{2}\Omega_{\rm{inf}}(f_{\Delta})e^{2N_{kin}-N_{m}}.$ (3.28)
@@
In the above, $f_{\Delta}$ is the frequency corresponding to the end of the matter domination at $t_{\Delta}$, and $N_{kin},N_{m}$ are the e-folds of kination and matter domination respectively. This must compared with (5.11), which is always larger by a factor $\left(\frac{\mu(t_{0})}{\mu{(t_{\Delta})}}\right)^{2}$, the ratio of the string tensions at the initial and final times.
@@ 3.2 Effect on the string tension
To close of this section, let us examine how the fundamental string tension depends on time in the examples we have considered. Let us remind the reader that this can be obtained by simply plugging in the time-dependence of the compactification volume [^30] And also of the dilaton, which here is constant. in (2.1). In the case of volume kination, one can combine Eqs (3.1) and (3.6) to obtain
$\mu_{\rm{kin}}(t)=\frac{g_{s}^{2}M_{P}^{2}}{4\pi\mathcal{V}_{0}^{\rm{kin}}}\frac{t_{0}^{\rm{kin}}}{t},$ (3.29)
@@
$\mu_{\rm{tr}}(t)=\frac{g_{s}^{2}M_{P}^{2}}{4\pi\mathcal{V}_{0}^{\rm{tr}}}\left(\frac{t_{0}^{\rm{tr}}}{t}\right)^{\frac{2}{\lambda}\sqrt{\frac{3}{2}}}.$ (3.30)
@@
$\mu^{LVS}_{\rm{tr}}(t)=\frac{g_{s}^{2}M_{P}^{2}}{4\pi\mathcal{V}_{0}^{\rm{tr}}}\left(\frac{t_{0}^{\rm{tr}}}{t}\right)^{\frac{2}{3}}.$ (3.31)